OpenStudy (rizags):

Disks in a tower of Hanoi are identified from top to bottom as 1, 2, 3, and so on. A computer can solve the puzzle by listing each disk moved, in order. For example, a 3-disk puzzle is solved as 1-2-1-3-1-2-1. If it takes a computer 1 second to display the 10-disc solution, how long would it take to display the solution to a 30 disc tower?

OpenStudy (rizags):

thanks

OpenStudy (aum):

Do you understand how this puzzle works?

OpenStudy (rizags):

i have played it before yes

OpenStudy (aum):

What needs to be done in the puzzle?

OpenStudy (rizags):

ok so for a 3 disk puzzle, there are 3 hoops and 3 disks on the first hoop. The smallest disk is on the top, followed by a medium size and a large disk. in one move, you can only move one disk to another hoop of your choice. you cannot move a disk on top of a larger one. I can link u a small flash version of the game on mathisfun http://www.mathsisfun.com/games/towerofhanoi.html

OpenStudy (aum):

Is there a formula for how many moves in a n-disc game?

OpenStudy (rizags):

i suppose this helped me http://mathworld.wolfram.com/TowerofHanoi.html

OpenStudy (aum):

Is this from a class textbook or from a puzzle book?

OpenStudy (rizags):

the question is from class but the link is from a textbook explanation

OpenStudy (rizags):

i just can't decipher the textbook

OpenStudy (rizags):

waiiiiiiiiiiiit

OpenStudy (rizags):

this formula works

OpenStudy (rizags):

h\[h_{n} = 2^{n} -1\]

OpenStudy (rizags):

where h subscript n is total moves and n is the number of disks

OpenStudy (aum):

Plug n = 10 into the formula and find the number of moves. Plug n = 30 into the formula and find the number of moves. Find the ratio. The display time will be in the same ratio.

OpenStudy (rizags):

that will produce 1023 and 1073741823, whose ratio is 1049601 THANKS!

OpenStudy (rizags):

wait i have one last problem that is baffling me

OpenStudy (aum):

Is 1049601 seconds the correct answer? Or do they want it in days, hours, minutes, seconds?

OpenStudy (rizags):

it doesnt matter i can solve it from this point

OpenStudy (rizags):

i just need help with one last

OpenStudy (aum):

go ahead.

OpenStudy (rizags):

Determine a, b, and c if the number at the end of each row in the chart shown below is the sum of the numbers in the row. Find a rule that relates the numbers at the bottom of each column to the entries in the column.

OpenStudy (rizags):

sorry for bad drawing

OpenStudy (aum):

1st row: b+c+a+c = a+b+2c = 54 --- (1) 2nd row: a+c+b+a = 2a+b+c = 50 --- (2) 3rd row: b+b+c+a = a+2b+c = 44 --- (3) Three equations and three unknowns. Solve for a, b and c.

OpenStudy (rizags):

so a+b+c=37 which means a=13 b=7 c=17

OpenStudy (rizags):

is that the whole thing? i dont understand the rule part

OpenStudy (rizags):

i mean, i understand it, im just having trouble locating the rule

OpenStudy (aum):

Redraw the table with a,b,c filled in. Drop the last column (the row totals column) in the new table. We can take a look at it and see if some pattern pops up.

OpenStudy (rizags):